### v1

@@ Abstract
This study investigates the molecular-level mechanism of Alternating Current Electrohydrodynamic (AC-EHD) flow in nanopores under high-frequency conditions, using molecular dynamics simulations. A gold-NaCl system with symmetric and asymmetric electrode configurations is used to analyze the flow patterns under high-frequency AC potentials. Our findings reveal localized heat generation near the electrode leading to a steep temperature gradient. An order parameter analysis explained that the heat generation is due to the periodic change in the alignment of water molecules under AC potential and that at these high frequencies the influence of Na+ and Cl- ions are negligible. The heat generation and temperature gradient are found to increase with the applied AC frequency. Three different electrode configurations were studied by varying the size and distance between the electrodes. A net directional flow develops in the asymmetric electrode structures. A possible mechanism for this is proposed by analyzing the flow patterns using velocity and temperature profiles, order parameters, streamline plots and mean square displacements. Different forces acting on the fluid were identified such as buoyancy-driven convection due to temperature gradient, electrothermal effects influenced by the temperature-dependent properties of water, and Maxwell stress due to the non-uniform electric field. Moreover, the asymmetric electrode structure created an imbalance in these forces and generated a net directional flow. These findings suggest the existence of a form of nanoscale AC-EHD flow that operates in a frequency regime above that of conventional electroosmotic and electrothermal mechanisms and that, unlike these mechanisms, occurs independently of ionic concentration. Thereby this work provides insights for optimizing AC-EHD flow in nanoscale systems where precise fluid manipulation is critical.
@@ I INTRODUCTION
Electrohydrodynamics (EHD) is an interdisciplinary science between electrical and fluid mechanics that studies the dynamics of electrically charged or polarizable liquids due to the action of electric fields. The domain of EHD has potential for substantial theoretical and practical application ranging from microfluidics and the enhancement of heat transfer to electrostatic precipitation [1, 2, 3, 4]. Interaction between charges or dipoles in a fluid medium either creates flow or amplifies flow patterns, thus enabling the alteration of fluid behavior without using mechanical components [5]. Alternating Current Electrohydrodynamic (AC-EHD) flow represents one branch of EHD that applies AC electric fields to induce fluid motion. Applying an AC potential prevents the electrolysis and electrode fouling problems, which are typical in DC systems. This makes it much more suited for use in delicate biochemical processes or wherever chemical purity is a concern. The AC-EHD method utilizes time-averaged forces, such as dielectrophoresis and electrothermal forces, which are quite different from those used in DC-EHD systems [6]. The use of AC-EHD is particularly notable in microscale applications where precise control over fluid mixing, heat transfer, and mass transport is essential. These capabilities are necessary for the development of technologies like lab-on-a-chip devices, precision cooling systems, and micro-reactors [7, 8, 9, 10]. The recent advancements in understanding the AC-EHD phenomena have led to the designing of complex systems that employ the advantages of fluid dynamics for better performance.
@@
Over the years, crucial technological advances have harnessed the potential of AC-EHD and electrokinetic forces for isolation and detection of exosomes [11, 12], DNA trapping and analysis [13, 14], protein biomarker analysis [15, 16], cancer cell isolation and analysis [17, 18], and manipulation of cells and microorganisms [19, 20]. These advances have enabled sensitive detection and precise manipulation of clinically relevant biomarkers from complex biological samples. Shiddiky et al. [21] introduced a technique called AC-EHD-induced nano shearing to improve the detection of molecular biomarkers in complex biological samples. This method uses an alternating electric field to displace molecules that are nonspecifically bound to solid surfaces. The force generated by the electric field acts within nanometers of the electrode surface, shearing off loosely bound molecules and reducing nonspecific adsorption by up to five times. Additionally, this technique enhances the specific capture of target proteins, leading to a 1000-fold increase in the detection of HER2 protein from human serum, with sensitivity down to 1 fg mL-1. This approach allows for controlled shear forces and fluid micro-mixing, improving sensor-target interactions and making it suitable for clinical applications and early disease diagnosis.
@@
Vaidyanathan et al. [22] have developed a microfluidic platform by using AC-EHD induced nano shearing for capturing and detection of highly specific exosomes which is crucial for cancer biomarker analysis. This method utilizes AC-EHD-induced electrical forces near the electrode surface to enhance the capture specificity and reduce nonspecific adsorption. The device effectively isolated exosomes from breast cancer samples and showed that the detection sensitivity is three times better than hydrodynamic flow assays. Their platform using asymmetric pairs of microelectrodes is a promising tool for the rapid quantification of exosomes, diagnosis of cancer other biological applications.
Researchers have employed various experimental and analytical methodologies to understand the mechanism behind AC-EHD flow, thereby optimizing the conditions for its effective implementation. Brown et al. [23] investigated the Rotating Wave Approximation (RWA) for two-level systems and derived simple analytical solutions for the physical observables of the system under the interaction of laser. They have extended the RWA to many-level systems that interact with continuous wave lasers and discussed the effects of diagonal dipole matrix elements. The generalized RWA (GRWA) is beneficial for solving numerical equations integrations of the time-dependent Schrödinger equation thereby improving the computational efficiency. This model serves as a foundation for the development of GRWA for many-level systems interacting with pulsed lasers.
@@
Ramos et al. [24] conducted a theoretical analysis of electro-osmotic pumping which is driven by a kilohertz-scale AC electric potential applied across pairs of asymmetric microelectrodes. A net fluid flow is generated due to the interaction between the oscillating electric field and the induced charges in the diffused double layer of the electrode. They numerically solve the electrical equation using the charge simulation method, and the computational efficiency in the calculation is enhanced by considering the periodicity of the system. Finally, they have identified optimal non-dimensional parameters for enhancing the pumping velocity and also fluid flow dependence on voltage and frequency. Their results are well aligned with the experimental data reported by Brown et al. [23].
@@
González et al. [25] studied electrothermal motion in aqueous solutions subject to megahertz-scale electric fields and temperature gradients. They have developed solutions for fluid flow above two types of microelectrode structures. The first structure is two coplanar electrodes with an AC potential difference which generates two-dimensional rolls, and the second structure is four planar coplanar electrodes with a four-phase AC signal, creating a rotating liquid whirl. The theoretical predictions are confirmed by the experiments under strong illumination. Their findings suggest that such electrothermal flows could enhance mixing in microsystems where the flow velocity is independent of the solution conductivity and the efficiency can be improved by tuning the AC frequency.
@@
Different mechanisms have been reported to explain AC-EHD flow across various frequency ranges. Among these, two distinct mechanisms are particularly prominent in the literature. The first mechanism involves applying an alternating potential difference across asymmetric electrodes, creating a nonuniform electric field. The normal component of the field will induce a charge in the diffuse double-layer and the tangential component will exert a force on this induced charge. Since the sign of the charge reverses with the field, the force has a nonzero time average which results in an electro-osmotic slip velocity with both oscillatory and time-averaged components. The time-averaged velocity predominates in the the hundreds to the kilohertz range frequencies causing an observed fluid flow. This setup with coplanar asymmetric electrodes produces a localized flow that generates a global flow in the direction of broken symmetry.[24, 26, 6, 27]
@@
The second mechanism is electrothermally induced fluid flow, in which the electric fields interact with temperature gradients within an aqueous solution, causing fluid motion. These temperature gradients can be generated by different external sources, such as strong illumination, or through Joule heating from the applied electric field. The gradients will create an inhomogeneity in the system due to the variation of electrical conductivity and permittivity of fluid. When an alternating current electric field is applied, it exerts forces on these inhomogeneities, inducing fluid motion. Depending on the electrode configuration, this mechanism can produce different flow patterns: two coplanar electrodes with an AC potential difference generate two-dimensional rolls, while four coplanar electrodes with a four-phase AC signal create a rotating whirl. This electrothermal motion is useful in microsystems for mixing analytes and liquids, which offers advantages such as efficiency and independence of flow velocity from solution conductivity [25, 28, 29]. Even though various theories exist in the literature explaining the underlying mechanism of AC-EHD flow, there remains a need for more accurate molecular-level explanations, especially at higher frequencies and nanoscale dimensions. Further research is essential to fully understand and optimize the flow for various applications.
@@
As systems reduce in size to the nanoscale, the mechanisms are likely to differ from those proposed previously. In this paper, we investigate the molecular-level mechanism of nanoscale AC-EHD flow with a gold-NaCl aqueous system using molecular dynamics simulations. An AC potential with a very high-frequency range was employed using the constant potential method to analyze flow patterns. Various electrode arrangements were tested to determine the optimum geometry for achieving maximum flow velocity. Additionally, a theoretical explanation for AC-EHD flow at the nanoscale was developed based on the observed results. The following sections provide a summary of the modeling approach, the simulation methodology employed, a discussion of the obtained results, and, finally, some concluding remarks.
@@ II METHODOLOGY
Classical molecular dynamics (MD) simulations were carried out to investigate the underlying mechanism of alternating current electrohydrodynamic flow. The simulation setup consists of an aqueous solution of NaCl confined between two gold nanoplates. The schematic depiction of one of the models (model 1) used in this study is shown in Fig. 1. The top Au plate is kept as flexible, where the Au atoms are tethered around their initial lattice positions using a weak harmonic spring potential. The bottom gold plate consists of both electrodes placed at a distance and the atoms were kept rigid. The dimensions of the Au sheets are Lx= Ly=5.5 nm and the height of the channel is Lz=7 nm. The length of both electrodes is D=1.375 nm and the gap between the electrodes is also D=1.375 nm for this particular model. There are other models are also investigated with different electrode sizes, and gaps between electrodes. Periodic boundary conditions were applied in the $x$ and $y$ directions, and confinement was in the $z$-direction.
@@
The water molecules were modeled using the extended simple point charge (SPC/E) [30, 31] water model. The long-range electrostatic forces were calculated by using the particle-particle-particle-mesh (PPPM) [32] solver with an accuracy of 1 $\times$ 10-6, and the water molecules were kept rigid by using the SHAKE [33] algorithm. To use the PPPM method, periodic boundary conditions in all directions are required. An Ewald summation technique with an extended volume ratio of 3.0 was used due to the confinement in the $z$ direction [34]. The interactions between particles are of the Lennard-Jones form and the parameters for the cross-terms are derived using the Lorentz-Berthelot mixing rules. The force field parameters for Au(100) were taken from Walsh et al. [35]. The non-polarisable NaCl with SPC/E water models were modeled using the parameters taken from Weerasinghe and Smith [36]. A velocity-Verlet algorithm was used to integrate the equations of motion for every particle, with a time step of 1 fs. All of the MD simulations were carried out using the LAMMPS (large-scale atomic/molecular massively parallel simulator) package [37], while the CPM was carried out using the ELECTRODE [38] package. Visual MD [39] is used for model rendering and visualization.
Initially, the system was equilibrated in a canonical (NVT) ensemble at a reference temperature of 300 K for a time period of 2.0 ns. In addition, the system stability was checked by simulating in a microcanonical (NVE) ensemble for another 2.0 ns. Once the system is properly equilibrated, a constant potential method was applied to the electrodes, and the top gold plate was thermostatted at 300 K for 2 ns. Finally, all the analyses were carried out during another 2 ns production simulation by maintaining the electrode potential and thermostat. A 4 V potential difference was applied to the electrodes using the constant potential method for both AC and DC. The AC frequency was varied from a few MHz to 100 GHz to understand the range of frequency to make any impact on the system at this timescale. Since the time scale used in this molecular dynamics simulation was in the nanosecond range, at least 1 GHz frequency is required to make a complete cycle of AC potential for 1 ns time. The cumulative heat removed from the system is calculated using the energy removed from the thermostat.
@@ III RESULTS AND DISCUSSIONS
The surface charge density of electrodes under different frequencies was investigated by changing the frequency from 1 MHz to 100 GHz. For the lower frequencies, the change in polarity during the time scale used in this study (nanoseconds) is not enough to have any observable influence on the system. Therefore we used very high frequencies – up to 100 GHz – to understand the molecular level mechanism of AC-EHD flow. Figure 2 shows the applied voltage and induced surface charge as a function of simulation time for a few selected frequencies. For the case of DC voltage, surface charge density starts from zero and reaches the maximum value of 4.0 V and remains constant throughout the simulation time. For a frequency of 100 MHz, the period is 10 ns so the voltage remains small during the 1 ns simulation time. For frequencies lower than this, it will take a very large simulation time to obtain the peak value of the voltage (first peak of sine wave). A 250 MHz frequency range is required to reach the peak voltage in 1 ns. From Fig. 2 we can see that a complete cycle is obtained when the frequency is 1 GHz. Therefore to study the effect of frequency on the system using molecular simulations in a viable timeframe, we need a very high frequency. From Fig 2 it is also evident that the response time of induced charge is very fast, as the surface charge density nearly follows the voltage curve even for higher frequencies.
@@
As mentioned in the methodology section, the AC potential is applied after the equilibration of the system. Once the constant potential is applied, the induced charge on the electrode changes its polarity according to the frequency of the AC potential applied. While analyzing the temperature of the system it was observed that the system temperature is increasing with time and reaches a steady state value after a particular time. Figure 3 shows the temperature of the fluid and heat removed by the thermostat for DC voltage and different AC voltage frequencies. In the case of DC voltage, the temperature of the fluid remains the same throughout the simulation period. However, it is observed that when the frequency of the AC is above 1 GHz, the temperature of the fluid increases and reaches a steady state. The temperature rise is proportional to the frequency of the applied AC potential. The increase in fluid temperature indicates that heat is generated in the system while applying AC potential, as would be expected. Figure 3b shows the heat removed by the thermostat for different frequencies. The heat removed by the thermostat remains constant for DC potential indicating zero heat generation after some time. When the AC potential is applied, the heat removed by the thermostat increases over time, indicating continuous heat generation that also depends on the frequency.
@@
To understand the pattern of heat generation in the system, the temperature profile and heat map of one of the systems are plotted in Fig 4. The temperature profile is determined by calculating local temperatures using the degrees of freedom correction method, which accounts for constraints in the molecular system[40, 41]. The temperature profile indicates that the temperature is maximum at the bottom plate, where the electrodes are placed, and gradually decreases towards the top plate. Therefore, it is understood that heat generation does not occur in the fluid as a whole; instead, heat is generated locally near the electrodes. In addition, the temperature contour plot shows that there are no heat generation spots along the electrode surface; instead, a uniform increase of temperature over the electrodes is visible. This could change if the electrodes increase in size.
@@
As mentioned in the methodology section, only the top plate is thermostatted and the bottom plate is kept rigid to simplify the application of the constant potential method. Therefore, to confirm the local heating near the electrodes, the effect of the thermostat should be analyzed. Figure 5 shows the effect of different thermostat settings on temperature rise and heat removal. For the first setting, where no thermostats are used, the temperature of the system continuously increases. This is because the heat generated in the system is not removed due to the absence of any thermostat and would represent the case of a system between thermally insulating plates. The second setting is where one thermostat is provided at the top plate, and the bottom plate is kept rigid, representing a top plate with high thermal conductivity. Finally, the third setting is used in which both top and bottom plates are thermostatted. Since the bottom plate is kept fixed due to the application of CPM, in order to do this an additional layer of flexible gold plate is introduced by overlapping the position of the fixed gold plate. The interactions between the fixed and flexible electrodes were turned off. The detailed methodology used for overlapping electrodes is available in our previous work [42]. This would represent the case where both plates have high thermal conductivity. Even though the rate of temperature rise is reduced when a thermostat is applied at the bottom plate, a temperature rise still occurs. From the data showing the heat removed by thermostat data in Fig 5b, it is clear that more heat is removed from the fluid when two thermostats are applied. This is the reason for a lower rate of temperature rise in the two thermostat systems.
@@
From the above analysis, it is clear that heat is being generated in the fluid nearer to the electrodes when a very high-frequency AC potential is applied. To understand the role of charged particles (Na+ and Cl- ions) in this phenomenon, the temperature rise of water and aqueous solution of NaCl is compared in Fig 6. Interestingly, the temperature rise is found to be similar for both systems, with and without NaCl. This suggests that the localized temperature rise is not caused by the movement of ions in the system under rapid changes in electrode polarity.
@@
To understand the mechanism behind the temperature rise, an orientational order parameter analysis of water molecules has been carried out [43]. The order parameter is defined as

### v2

@@ Abstract
This study investigates the molecular-level mechanism of Alternating Current Electrohydrodynamic (AC-EHD) flow in nanopores under high-frequency conditions, using molecular dynamics simulations. A gold-NaCl system with symmetric and asymmetric electrode configurations is used to analyze the flow patterns under high-frequency AC potentials. Our findings reveal localized heat generation near the electrode leading to a steep temperature gradient. An order parameter analysis indicates that the heat generation is due to the periodic change in the alignment of water molecules under AC potentials. At these high frequencies the influence of Na+ and Cl- ions are negligible. The heat generation and temperature gradient are found to increase with the applied AC frequency. Three different electrode configurations were studied by varying the size and distance between the electrodes. A net directional flow develops in the asymmetric electrode structures. A possible mechanism for this is proposed by analyzing the flow patterns using velocity and temperature profiles, order parameters, streamline plots and mean square displacements. Different effects on the fluid were identified including those associated with temperature gradients, temperature-dependent fluid properties, and non-uniform electric fields. The asymmetric electrode structure created an imbalance in these effects and generated a net directional flow. These findings suggest the existence of a form of nanoscale AC-EHD flow that operates in a frequency regime above that of conventional electroosmotic and electrothermal mechanisms and that, unlike these mechanisms, occurs independently of ionic concentration. Thereby this work provides insights for optimizing AC-EHD flow in nanoscale systems where precise fluid manipulation is critical.
@@ I INTRODUCTION
Electrohydrodynamics (EHD) is an interdisciplinary science between electrical and fluid mechanics that studies the dynamics of electrically charged or polarizable liquids due to the action of electric fields. The domain of EHD has potential for substantial theoretical and practical application ranging from microfluidics and the enhancement of heat transfer to electrostatic precipitation [jones1995electromechanics, melcher1969electrohydrodynamics, velev2006chip, bazant2004induced]. Interaction between charges or dipoles in a fluid medium either creates flow or amplifies flow patterns, thus enabling the alteration of fluid behavior without using mechanical components [castellanos2014electrohydrodynamics]. Alternating Current Electrohydrodynamic (AC-EHD) flow represents one branch of EHD that applies AC electric fields to induce fluid motion. Applying an AC potential prevents the electrolysis and electrode fouling problems, which are typical in DC systems. This makes it much more suited for use in delicate biochemical processes or wherever chemical purity is a concern. The AC-EHD method utilizes time-averaged forces, such as dielectrophoresis and electrothermal forces, which are quite different from those used in DC-EHD systems [ramos1999ac]. The use of AC-EHD is particularly notable in microscale applications where precise control over fluid mixing, heat transfer, and mass transport is essential. These capabilities are necessary for the development of technologies like lab-on-a-chip devices, precision cooling systems, and micro-reactors [ramos1998ac, nguyen2006fundamentals, silva20223d, ronkainen2010electrochemical]. The recent advancements in understanding the AC-EHD phenomena have led to have led to the design of complex systems that employ the advantages of their fluid dynamics for better performance.
@@
Over the years, crucial technological advances have harnessed the potential of AC-EHD and electrokinetic forces for isolation and detection of exosomes [wei2013detection, boriachek2018biological], DNA trapping and analysis [basuray2009shear, kumemura2011single], protein biomarker analysis [gong2010label, vaidyanathan2015multiplexed], cancer cell isolation and analysis [gupta2012apostream, dey2016electric], and manipulation of cells and microorganisms [kwon2018microfluidic, ren2018bioparticle]. These advances have enabled sensitive detection and precise manipulation of clinically relevant biomarkers from complex biological samples. Shiddiky et al. [shiddiky2014molecular] introduced a technique called AC-EHD-induced nano shearing to improve the detection of molecular biomarkers in complex biological samples. This method uses an alternating electric field to displace molecules that are nonspecifically bound to solid surfaces. The force generated by the electric field acts within nanometers of the electrode surface, shearing off loosely bound molecules and reducing nonspecific adsorption by up to five times. Additionally, this technique enhances the specific capture of target proteins, leading to a 1000-fold increase in the detection of HER2 protein from human serum, with sensitivity down to 1 fg mL-1. This approach allows for controlled shear forces and fluid micro-mixing, improving sensor-target interactions and making it suitable for clinical applications and early disease diagnosis.
@@
Vaidyanathan et al. [vaidyanathan2014detecting] have developed a microfluidic platform by using AC-EHD induced nano shearing for capturing and detection of highly specific exosomes which is crucial for cancer biomarker analysis. This method utilizes AC-EHD-induced electrical forces near the electrode surface to enhance the capture specificity and reduce nonspecific adsorption. The device effectively isolated exosomes from breast cancer samples and showed that the detection sensitivity is three times better than hydrodynamic flow assays. Their platform using asymmetric pairs of microelectrodes is a promising tool for the rapid quantification of exosomes, diagnosis of cancer and other biological applications.
Researchers have employed various experimental and analytical methodologies to understand the mechanism behind AC-EHD flow, thereby optimizing the conditions for its effective implementation. Brown et al. [brown2000pumping] consider asymmetric electrodes where flow is induced by a gradient in the potential parallel to their surface. The voltages required to observe this are relatively small, and the dependence of the flow on the potential and frequency is studied in this work. A model is proposed that is able to qualitatively predict the observed fluid velocities, and factors that need to be considered to improve the model are described.
@@
Ramos et al. [ramos2003pumping] conducted a theoretical analysis of electro-osmotic pumping which is driven by a kilohertz-scale AC electric potential applied across pairs of asymmetric microelectrodes. A net fluid flow is generated due to the interaction between the oscillating electric field and the induced charges in the diffused double layer of the electrode. They numerically solve the electrical equation using the charge simulation method, and the computational efficiency in the calculation is enhanced by considering the periodicity of the system. Optimal non-dimensional parameters are identified for enhancing the pumping velocity and the dependence of the fluid flow on voltage and frequency is determined. Their results are well aligned with the experimental data reported by Brown et al. [brown2000pumping].
@@
González et al. [gonzalez2006electrothermal] studied electrothermal motion in aqueous solutions subject to both megahertz-scale frequency alternating electric fields and temperature gradients produced by illumination of the system. They have developed solutions for fluid flow using two types of microelectrode structures. The first structure is two coplanar electrodes with an AC potential difference which generates two-dimensional rolls, and the second structure is four planar coplanar electrodes with a four-phase AC signal, creating a rotating liquid whirl. The theoretical predictions are confirmed by the experiments under strong illumination. Their findings suggest that such electrothermal flows could enhance mixing in microsystems where the flow velocity is independent of the solution conductivity and the efficiency can be improved by tuning the AC frequency.
@@
Different mechanisms have been reported to explain AC-EHD flow across various frequency ranges. Among these, two distinct mechanisms are particularly prominent in the literature. The first mechanism involves applying an alternating potential difference across asymmetric electrodes, creating a nonuniform electric field. The normal component of the field will induce a charge in the diffuse double-layer and the tangential component will exert a force on this induced charge. Since the sign of the charge reverses with the field, the force has a nonzero time average which results in an electro-osmotic slip velocity with both oscillatory and time-averaged components. The time-averaged velocity predominates in the the hundreds to the kilohertz range frequencies causing an observed fluid flow. This setup with coplanar asymmetric electrodes produces a localized flow that generates a global flow in the direction of broken symmetry.[ramos2003pumping, ajdari1996generation, ramos1999ac, green2002fluid]
@@
The second mechanism is electrothermally induced fluid flow, in which the electric fields interact with temperature gradients within an aqueous solution, causing fluid motion. These temperature gradients can be generated by different external sources, such as strong illumination, or through Joule heating from the applied electric field. The gradients will create an inhomogeneity in the system due to the variation of electrical conductivity and permittivity of fluid. When an alternating current electric field is applied, the forces induced lead to fluid motion. Depending on the electrode configuration, this mechanism can produce different flow patterns: two coplanar electrodes with an AC potential difference generate two-dimensional rolls, while four coplanar electrodes with a four-phase AC signal create a rotating whirl. This electrothermal motion is useful in microsystems for mixing analytes and liquids, which offers advantages such as efficiency and independence of flow velocity from solution conductivity [gonzalez2006electrothermal, green2000electric, green2001electrothermally]. Even though various theories exist in the literature explaining the underlying mechanism of AC-EHD flow, there remains a need for more accurate molecular-level explanations, especially at higher frequencies and nanoscale dimensions. Further research is essential to fully understand and optimize the flow for various applications. Recent numerical studies have further advanced the understanding of AC-driven electrohydrodynamic phenomena at micro- and mesoscopic scales, highlighting the roles of AC frequency, interfacial polarization, and electric-field nonuniformity in governing flow and particle dynamics. In particular, Tao et al. demonstrated frequency-dependent electrohydrodynamic interactions and Maxwell-stress-mediated transport using fully coupled multiphysics simulations [tao2025many, tao2025alternating]. These studies provide important context for AC-EHD behavior at larger length scales and motivate the need for molecular-level investigation under extreme nanoscale confinement.
Experimental studies have demonstrated that AC-EHD excitation can generate localized near-surface fluid motion over asymmetric electrodes, often referred to as “nanoshearing”, even though the overall device dimensions remain at the microfluidic scale. In particular, Shiddiky et al.  [shiddiky2014molecular] and Vaidyanathan et al. [vaidyanathan2014detecting] reported AC-EHD-induced near-electrode flow confined to within a few nanometers of the electrode surface under kHz–MHz excitation. While these experiments confirm the physical existence of AC-EHD-driven near-surface flow, controlled measurements under extreme nanoscale confinement and GHz-frequency excitation are not currently available, motivating molecular-level investigation.
@@
As systems reduce in size to the nanoscale, the mechanisms are likely to differ from those proposed previously. In this paper, we investigate the molecular-level mechanism of nanoscale AC-EHD flow with a gold-NaCl aqueous system using molecular dynamics simulations. Molecular dynamics is used to capture molecular-scale and interfacial mechanisms that govern electrohydrodynamic flow under extreme nanoscale confinement, which are not explicitly resolved in mesoscopic approaches. An AC potential with a range of very high-frequency range was employedfrequencies is applied to the systems using the constant potential method to analyze (CPM) and the flow patterns. are analyzed. Various electrode arrangements wereare tested to determine how the optimum geometry for achievingcan be tuned to achieve maximum flow velocity. Additionally, a theoretical explanation for AC-EHD flow at the nanoscale wasis developed based on the observed results. The following sections provide a summary of the modeling approach, the simulation methodology employed, a discussion of the obtained results, and, finally, some concluding remarks.
@@ II METHODOLOGY
Classical molecular dynamics (MD) simulations were carried out to investigate the underlying mechanism of alternating current electrohydrodynamic flow. The simulation setup consists of an aqueous solution of NaCl confined between two gold nanoplates. The schematic depiction of one of the models (model 1) used in this study is shown in Fig. 1. The atoms in the top Au plate are flexible, with the Au atoms tethered around their initial lattice positions using a weak harmonic spring potential. The bottom gold plate contains electrodes placed at a fixed distance with rigid atoms. The dimensions of the Au sheets are Lx = Ly = 5.5 nm and the height of the channel is Lz = 7 nm. The length of both electrodes is D = 1.375 nm and the gap between the electrodes is also D = 1.375 nm for the model shown. Other models are also investigated with different electrode sizes and gaps between the electrodes. Periodic boundary conditions were applied in the $x$ and $y$ directions, and confinement was in the $z$-direction
@@
The water molecules were modeled using the extended simple point charge (SPC/E) [berendsen1987missing, wu2006flexible] water model. The long-range electrostatic forces were calculated by using the particle-particle-particle-mesh (PPPM) [hockney2021computer] solver with an accuracy of 1 $\times$ 10-6, and the water molecules were kept rigid by using the SHAKE [ryckaert1977numerical] algorithm. To use the PPPM method, periodic boundary conditions in all directions are required. An Ewald summation technique with an extended volume ratio of 3.0 was used due to the confinement in the $z$ direction [yeh1999ewald]. The non-Coulombic interactions between particles are of the Lennard-Jones form and the parameters for the cross-terms are derived using the Lorentz-Berthelot mixing rules. The force field parameters for Au(100) were taken from Wright et al. [wright2013golp]. The non-polarizable NaCl with SPC/E water models were modeled using the parameters taken from Weerasinghe and Smith [weerasinghe2003kirkwood]. A velocity-Verlet algorithm was used to integrate the equations of motion for every particle, with a time step of 1 fs. All of the MD simulations were carried out using the LAMMPS (large-scale atomic/molecular massively parallel simulator) package [plimpton1995fast], while the CPM was carried out using the ELECTRODE [siepmann1995influence] package. Visual MD [humphrey1996vmd] is used for model rendering and visualization.
Initially, the system was equilibrated in a canonical (NVT) ensemble at a reference temperature of 300 K for a time period of 2.0 ns. In addition, the system stability was checked by simulating the system in a microcanonical (NVE) ensemble for another 2.0 ns. Once the system was properly equilibrated, a constant potential method was applied to the electrodes, and the top gold plate was thermostatted at 300 K for 2 ns. Finally, all the analyses were carried out during another 2 ns production simulation by maintaining the electrode potential and thermostat. A 4 V potential difference was applied to the electrodes using the constant potential method for both AC and DC. The AC frequency was varied from a few MHz to 100 GHz to understand the range of frequency required to make any impact on the system at this timescale. Since the time scale used in this molecular dynamics simulation was in the nanosecond range, at least 1 GHz frequency is required to make a complete cycle of AC potential for 1 ns time. The cumulative heat removed from the system is calculated using the energy removed from the thermostat.
@@ III RESULTS AND DISCUSSIONS
The surface charge density of the electrodes with different frequencies was investigated for frequencies from 1 MHz to 100 GHz. For the lower frequencies, the change in polarity during the time scale used in this study (nanoseconds) is not enough to have any observable influence on the system. Therefore we used very high frequencies – up to 100 GHz – to understand the molecular level mechanism of AC-EHD flow. Figure 2 shows the applied voltage and induced surface charge as a function of simulation time for a few selected frequencies. For the case of DC voltage, surface charge density starts from zero and reaches the maximum value of 4.0 V and remains constant throughout the simulation time. For a frequency of 100 MHz, the period is 10 ns so the voltage remains small during the 1 ns simulation time. For frequencies lower than this, it will take a very large simulation time to obtain the peak value of the voltage (first peak of sine wave). A 250 MHz frequency range is required to reach the peak voltage in 1 ns. From Fig. 2 we can see that a complete cycle is obtained when the frequency is 1 GHz. Therefore to study the effect of frequency on the system using molecular simulations in a viable timeframe, we need a very high frequency. From Fig 2 it is also evident that the response time of induced charge is very fast, as the surface charge density nearly follows the voltage curve even for higher frequencies.
@@
As mentioned in the methodology section, the AC potential is applied after the equilibration of the system. Once the constant potential is applied, the induced charge on the electrode changes its polarity according to the frequency of the AC potential applied. While analyzing the temperature of the system it was observed that the system temperature is increasing with time and reaches a steady state value after a particular time. Figure 3 shows the temperature of the fluid and heat removed by the thermostat for DC voltage and different AC voltage frequencies. In the case of DC voltage, the temperature of the fluid remains the same throughout the simulation period. However, it is observed that when the frequency of the AC is approximately 1 GHz, the temperature of the fluid can be observed to increase and reach a steady state. The temperature rise is proportional to the frequency of the applied AC potential. The increase in fluid temperature indicates that heat is generated in the system while applying AC potential, as would be expected. Figure 3b shows the heat removed by the thermostat for different frequencies. The heat removed by the thermostat remains constant for DC potential indicating zero heat generation after some time. When the AC potential is applied, the heat removed by the thermostat increases over time, indicating continuous heat generation that also depends on the frequency. We note that the observed heating is spatially localized near the electrodes, present even in low-flow configurations, and results in steady-state temperatures roughly proportional to the applied AC frequency (3(a)). Thus, while our molecular dynamics simulations should include most physically relevant energy dissipation mechanisms, including viscous dissipation due to fluid motion, the observed heating is directly attributable to the AC frequency imposed rather than resulting predominantly from viscous dissipation.
@@
To understand the pattern of heat generation in the system, the temperature profile and heat map of one of the systems are plotted in Fig 4. The temperature profile is determined by calculating local temperatures using the degrees of freedom correction method, which accounts for constraints in the molecular system[sanderson2024local, sanderson5187317dofulator]. The temperature profile indicates that the temperature is maximum at the bottom plate, where the electrodes are placed, and gradually decreases towards the top plate. Therefore, it is understood that heat generation does not occur in the fluid as a whole; instead, heat is generated locally near the electrodes. In addition, the temperature contour plot shows that there are no heat generation spots along the electrode surface; instead, a uniform increase of temperature over the electrodes is visible. This could change if the electrodes increase in size.
@@
As mentioned in the methodology section, for the simulations presented so far only the top plate is thermostatted and the bottom plate is kept rigid to simplify the application of the constant potential method. Therefore, to confirm the local heating near the electrodes, the effect of the thermostat should be analyzed. Figure 5 shows the effect of different thermostat settings on temperature rise and heat removal. For the first setting, where no thermostats are used, the temperature of the system continuously increases. This is because the heat generated in the system is not removed due to the absence of any thermostat and would represent the case of a system between thermally insulating plates. The second setting is where one thermostat is provided at the top plate, and the bottom plate is kept rigid, representing a top plate with high thermal conductivity. Finally, the third setting is used in which both top and bottom plates are thermostatted. Since the bottom plate is usually kept fixed due to theenable efficient application of the CPM, in order to do this an additional layer of flexible gold plate sites is introduced which overlap the the position of the fixed gold plate sites. There are no interactions between the fixed and flexible electrodes sites. The detailed methodology used for overlapping electrodes is available in our previous work [alosious2024interfacial]. This would represent the case where both plates have high thermal conductivity. Even though the rate of temperature rise is reduced when a thermostat is applied to the bottom plate, a temperature rise still occurs. From the data showing the heat removed by thermostat data in Fig 5b, it is clear that more heat is removed from the fluid when two thermostats are applied. This is the reason for a lower rate of temperature rise in the two thermostat systems.
@@
From the above analysis, it is clear that heat is being generated in the fluid nearer to the electrodes when a very high-frequency AC potential is applied. To understand the role of charged particles (Na+ and Cl- ions) in this phenomenon, the temperature rise of water and aqueous solution of NaCl is compared in Fig 6. Interestingly, the temperature rise is found to be similar for both systems, with and without NaCl. This suggests that the localized temperature rise is not caused by the movement of ions in the system under rapid changes in electrode polarity. A simple order-of-magnitude diffusion-timescale estimate provides additional physical insight: the characteristic ionic diffusion length during one AC cycle scales as $\ell_{d}\sim\sqrt{2D/f}$, where $D$ is the ionic diffusivity and $f$ is the AC frequency. For typical ionic diffusivities in water ($D\sim 10^{-9}\,{\mathrm{m}^{2}\,\mathrm{s}^{-1}}$), ions cannot diffuse across the 7 nm channel within a single cycle above tens of MHz, and at GHz frequencies the per cycle diffusion distance is only on the order of a nanometer. This supports the reduced role of ion-mediated mechanisms at the high frequencies considered here.
@@
To understand the mechanism behind the temperature rise, an orientational order parameter analysis of water molecules has been carried out [de1993physics]. The order parameter is defined as